# On the Similarities of Embeddings in Contrastive Learning

### v1

@@ 4 Similarities of Embedding Pairs
The similarity between embeddings of a positive pair $({\bm{x}},{\bm{y}})$ is defined as
@@
dubbed as the positive-pair similarity for the encoder $f$. Similarly, the similarity between embeddings of a negative pair $({\bm{x}},{\bm{y}})$ is defined as
@@ 5.1 Full-Batch Contrastive Learning
Recall that we consider various contrastive loss functions which can be categorized into two parts: the InfoNCE-based contrastive loss in Def. 3.1 and the independently additive contrastive loss in Def. 3.2. Our first main result below provides the similarities of positive/negative pairs for the two types of contrastive losses, when the encoder is full-batch trained.
@@
Suppose $d\geq n-1$. Let the contrastive loss function $\mathcal{L}\!\left({\bm{U}},\!{\bm{V}}\right)$ be one of the following forms:
@@
Then, in the full-batch training, the embedding similarities for the optimal encoder $f^{\star}$ in (1) are
@@
According to Theorem 5.1, all positive pairs achieve the perfect alignment, while all negative pairs are uniformly separated with the cosine similarity of $-\frac{1}{n-1}$. This is a generalized version of previous studies (Lu & Steinerberger, 2022; Cho et al., 2024; Lee et al., 2024; Koromilas et al., 2024), where we extend in two directions. First, the CL loss formulations (Def. 3.1 and Def. 3.2) we considered are a much broader class of losses compared with existing work. Second, we consider the randomness of $n$ embedding pairs in the optimization, where this randomness is introduced through augmentations.
Now we analyze the behavior of embeddings for arbitrary encoder $f$, that is not necessarily in the optimal status $f^{\star}$. The below result provides the relationship between the positive-pair similarity and the negative-pair similarity.
@@
Theorem 5.2 highlights the relationship between the expectation of positive-pair similarities, $\mathbb{E}\left[{\bm{s}}\left(f;\hat{p}_{\operatorname{pos}}\right)\right]$, and that of negative-pair similarities, $\mathbb{E}\left[{\bm{s}}\left(f;\hat{p}_{\operatorname{neg}}\right)\right]$. When the average of negative-pair similarities drops below $-\frac{1}{n-1}$, the positive pairs cannot be fully aligned, which is not desired. We call such phenomenon as the excessive separation of negative pairs in full-batch CL, since the average of negative-pair similarities drops below $-\frac{1}{n-1}$ when negative pairs are more separated compared with the optimal status specified in Theorem 5.1. This issue may arise when certain loss functions are used, as shown in the following theorem.
@@
Suppose that $d\geq n$. Consider the contrastive loss function $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ in the form of $\mathcal{L}_{\operatorname{ind-add}}\left({\bm{U}},{\bm{V}}\right)\;$ in Def. 3.2, where $(c_{1},c_{2})=(1,0)$ and
@@
Then, under the full-batch training setup, the embedding similarities for the optimal encoder $f^{\star}$ in (1) satisfy
@@
The inequality condition on $\phi$ and $\psi$ in (8) explains why the loss $\mathcal{L}_{\operatorname{ind-add}}({\bm{U}},{\bm{V}})$ in Def. 3.2 causes the excessive separation of negative pairs. Specifically, $\mathcal{L}_{\operatorname{ind-add}}({\bm{U}},{\bm{V}})$ is formulated as the sum of $\psi(\cdot)$ over negative pairs and $-\phi(\cdot)$ over positive pairs. Therefore, $\psi^{\prime}(-\frac{1}{n-1})$ indicates how much the loss decreases if the negative-pair similarity falls below $-\frac{1}{n-1}$, while $\phi^{\prime}(1)$ measures how much the loss increases if the positive-pair similarity drops below $1$. When the condition in (8) is satisfied, ignoring the scaling factor, the loss reduction from separating negative pairs beyond the similarity threshold of $-\frac{1}{n-1}$ can be greater than the loss increase from reducing positive-pair similarity below $1$. As a result, the optimization process favors pushing negative pairs even further apart, leading to the excessive separation. We provide a specific loss function where this issue arises:
@@
Addressing Excessive Separation in Full-Batch CL.
A natural question that arises is whether the excessive separation of negative pairs in full-batch CL can be mitigated. According to Theorem 5.1 and Theorem 5.3, this issue depends on the specific form of the contrastive loss. In particular, under the independently additive loss function $\mathcal{L}_{\operatorname{ind-add}}({\bm{U}},{\bm{V}})$, the optimal embeddings do not suffer from excessive separation when $c_{2}=1$ (i.e., case (ii) of Theorem 5.1), whereas the issue arises when $c_{2}=0$ and condition (8) holds.
This observation suggests two potential solutions to the excessive separation problem. The first is to set $c_{2}=1$ in the loss function, effectively incorporating within-view negative pairs. The second is to tune hyperparameters such that condition (8) does not hold, i.e., case (iii) in Theorem 5.1. While both approaches are theoretically valid, the former offers a more principled and practical remedy. In contrast, the latter lacks clear guidance for selecting suitable hyperparameters and may incur significant computational overhead. Therefore, we advocate including within-view negative pairs as a straightforward and effective strategy to avoid excessive separation in full-batch CL.
@@ 5.2 Mini-Batch Contrastive Learning
In Sec. 5.1, we show that when a proper contrastive loss function is chosen for full-batch training, the learned embedding satisfies the following behavior: all negative pairs have the cosine similarity of $-\frac{1}{n-1}$ and all positive pairs are perfectly aligned. What about the practical scenarios when we use mini-batches for training? Suppose $n$ training samples are partitioned into $b$ mini-batches where each batch contains $m:=n/b$ samples. Consider training the embeddings by under the fixed mini-batch configuration, where $k$-th mini-batch contains the samples with indices in ${\bm{I}}_{k}=\{m(k-1)+1,m(k-1)+2,\cdots,mk\}$. Under this scenario, the below theorem analyzes the behavior of embeddings learned by mini-batch training.
@@
Suppose $d\geq m-1$. Let the contrastive loss function $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ is one of the forms in Theorem 5.1. Define $f^{\star}_{\operatorname{batch}}$ as the optimal encoder that minimizes the fixed mini-batch loss, given by
@@
According to Theorem 5.5, the embeddings learned by mini-batch training have the following behaviors. First, the positive-pair similarity is equal to 1, i.e., all positive pairs are fully aligned. Second, the expectation of negative-pair similarities is equal to $-\frac{1}{n-1}$, which happens for the full-batch training as well. Third, unlike full-batch training, the negative-pair similarity is not uniform across the pairs, i.e., the variance is positive when the mini-batch size $m$ is strictly less than the sample size $n$. Thus, the effect of using mini-batch (compared with using full-batch) is in the increased variance of negative-pair similarities. Throughout the paper, we call such phenomenon as the excessive separation of negative pairs in mini-batch CL.
Figure 2 visualizes the effect of using mini-batches, compared with the full-batch scenario, when $n=4$ and $m=2$. For the full-batch training, shown in (c) of Figure 2, the variance of negative-pair similarities is zero, indicating that all negative pairs are equi-distant. In contrast, (a) and (b) of Figure 2 illustrate the embeddings learned by the mini-batch training, where the fixed mini-batches are specified as $\left({\bm{U}}_{[1:2]},{\bm{V}}_{[1:2]}\right)$ and $\left({\bm{U}}_{[3:4]},{\bm{V}}_{[3:4]}\right)$. For both (a) and (b), the variance of negative-pair similarities is positive, where (a) represents the case that achieves the upper bound of the variance in (11), and (b) corresponds to the case that achieves the lowest variance.
@@
Note that the variance of negative-pair similarities in (11) depends on the batch size $m$. For example, in the full-batch case where $m=n$, the upper bound in  (11) is zero, which is consistent with Theorem 5.1. One can confirm that the upper and lower bounds on the variance is a monotonically decreasing function of $m$, which implies that smaller batch sizes inherently exacerbates the excessive separation of negative pairs in mini-batch CL.
The below theorem analyzes the effect of batch size on the training dynamics, when a popular CL loss function is used:
@@
Addressing Excessive Separation in Mini-Batch CL.
Recall that the effect of using mini-batch in CL is in the increased variance of negative-pair similarities as shown in Theorem 5.5, and such phenomenon is called the excessive separation of negative pairs. To mitigate such issue, we introduce an auxiliary loss term $\mathcal{L}_{\operatorname{VRNS}}$ that reduces the variance of negative-pair similarities:
@@
as the auxiliary loss term for reducing the variance of negative-pair similarities.
One can combine arbitrary conventional mini-batch loss function $\mathcal{L}\left({\bm{U}}_{[m]},{\bm{V}}_{[m]}\right)$ with the proposed auxiliary loss to get the modified loss, given by
@@
where $\lambda>0$ is a hyperparameter. By including the proposed term into the mini-batch loss function, we encourage all negative-pair similarities to be close to $-\frac{1}{n-1}$, the ideal value achieved in the full-batch setting in Theorem 5.1. As a result, the proposed loss controls the variance of negative-pair similarities in mini-batch CL.
@@ 6 Empirical Validation
In this section, we empirically validate the impact of our theoretical results discussed in Sec. 5, especially for the practical scenarios of mini-batch training. First, we empirically observe that the excessive separation of negative pairs (proven in Theorem 5.5) actually happens in experiments on benchmark datasets. Second, we empirically confirm that such excessive separation issue can be mitigated by using the proposed loss term in Def. 5.7 which reduces the variance of the negative-pair similarities. Third, we observe such variance reduction improves the quality of learned representations in various real-world experiments.
@@ 6.1 Excessive Separation of Negative Pairs
To evaluate the effectiveness of our proposed auxiliary loss, we train five models for each batch size by minimizing the SimCLR loss combined with the auxiliary loss $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$ in Def. 5.7 with the hyperparameter of $\lambda=30$. The variances of negative-pair similarities from these additional models are shown in the last column of Table 1, and are consistently reduced across all batch sizes. This indicates that our proposed loss effectively mitigates excessive separation of negative pairs in mini-batch training.
@@ 6.2 Effect of Variance Reduction on Performance
We further investigate whether reducing the variance of negative-pair similarities actually improves the quality of learned representations in terms of the downstream performances.
@@
We pretrain models on CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009) using various contrastive loss functions that follow the formulation in Def. 3.1, including SimCLR, DCL, and DHEL. For all methods, we compare models trained with and without incorporating the auxiliary loss $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$ introduced in Def. 5.7. The hyperparameter $\lambda$ for the proposed loss is tuned over $\{0.1,0.3,1,3,10,30,100\}$. Unless otherwise specified, the other settings follow those in Sec. 6.1.
@@
Figure 3: Classification accuracy on CIFAR datasets. Models are trained by minimizing the SimCLR loss with and without the auxiliary loss term proposed in Def. 5.7, using various temperature parameters in the SimCLR loss.
The quality of the representations learned through CL is known to be sensitive to the choice of the temperature parameter in contrastive losses, as it influences the distribution of similarities among embeddings (Wang & Liu, 2021). To investigate this sensitivity, we train models using the SimCLR loss with temperature values ranging from 0.07 to 2.00. We compare the standard SimCLR loss against our proposed variant, which incorporates the auxiliary loss term introduced in Def. 5.7 to reduce the variance of negative-pair similarities. As shown in Figure 3, incorporating the proposed term leads to a consistently higher and more stable classification accuracy across all temperature settings, alleviating the need for careful temperature tuning.
Figure 4: Effect of the auxiliary loss term proposed in Def. 5.7 on top-1 classification accuracy when combined with various baseline methods (SimCLR, DCL, and DHEL) on CIFAR datasets. The gray bars highlight the performance gains achieved by incorporating the proposed term across various batch sizes. The proposed auxiliary loss term consistently improves the model performance.
We further evaluate the effect of the auxiliary term in Def. 5.7, when applied to existing CL methods including DCL and DHEL. Experiments are conducted on the CIFAR datasets with various batch sizes. As shown in Figure 4, incorporating the proposed loss term leads to improved classification accuracy, with the effect being more pronounced at smaller batch sizes. Additional results on the ImageNet dataset are presented in Appendix E.

### v2

@@ 4 Similarities of Embedding Pairs
The similarity of embeddings of a positive pair is defined as
@@
dubbed as the positive-pair similarity for the encoder $f$. Similarly, the similarity of embeddings of a negative pair is defined as
@@ 5.1 Full-Batch Contrastive Learning
Recall that we consider various contrastive losses which can be categorized into two parts: the InfoNCE-based contrastive loss in Def. 3.1 and the independently additive contrastive loss in Def. 3.2. Our first main result below provides the similarities of positive/negative pairs for the two types of contrastive losses, when the encoder is full-batch trained.
@@
Suppose $d\geq n-1$. Let the contrastive loss $\mathcal{L}\!\left({\bm{U}},\!{\bm{V}}\right)$ be one of the following forms:
@@
Then, in full-batch settings, the embedding similarities for the optimal encoder $f^{\star}$ in (1) are
@@
According to Theorem 5.1, all positive pairs achieve the perfect alignment, while all negative pairs are uniformly separated with the cosine similarity of $-\frac{1}{n-1}$. This is a generalized version of previous studies (Lu & Steinerberger, 2022; Cho et al., 2024; Lee et al., 2024; Koromilas et al., 2024), where we extend in two directions. First, the CL loss formulations (Def. 3.1 and Def. 3.2) we considered are a much broader class of losses compared with existing work. Second, we consider the randomness of $n$ embedding pairs in the optimization, where this randomness is introduced through augmentations. Now we analyze the behavior of embeddings for arbitrary encoder $f$, that is not necessarily in the optimal status $f^{\star}$. The below result provides the relationship between the positive-pair similarity and the negative-pair similarity.
@@
Theorem 5.2 highlights the relationship between the expectation of positive-pair similarities, $\mathbb{E}\left[{\bm{s}}\left(f;\hat{p}_{\operatorname{pos}}\right)\right]$, and that of negative-pair similarities, $\mathbb{E}\left[{\bm{s}}\left(f;\hat{p}_{\operatorname{neg}}\right)\right]$. When the average of negative-pair similarities drops below $-\frac{1}{n-1}$, the positive pairs cannot be fully aligned, which is not desired. We call such phenomenon as the excessive separation of negative pairs in full-batch CL, since the average of negative-pair similarities drops below $-\frac{1}{n-1}$ when negative pairs are more separated compared with the optimal status specified in Theorem 5.1. This issue may arise when certain losses are used, as shown in the following theorem.
@@
Suppose that $d\geq n$. Consider the contrastive loss $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ in the form of $\mathcal{L}_{\operatorname{ind-add}}\left({\bm{U}},{\bm{V}}\right)\;$ in Def. 3.2, where $(c_{1},c_{2})=(1,0)$ and
@@
Then, under full-batch settings, the embedding similarities for the optimal encoder $f^{\star}$ in (1) satisfy
@@
The inequality condition on $\phi$ and $\psi$ in (8) explains why the loss $\mathcal{L}_{\operatorname{ind-add}}({\bm{U}},{\bm{V}})$ in Def. 3.2 causes the excessive separation of negative pairs. Specifically, $\mathcal{L}_{\operatorname{ind-add}}({\bm{U}},{\bm{V}})$ is formulated as the sum of $\psi(\cdot)$ over negative pairs and $-\phi(\cdot)$ over positive pairs. Therefore, $\psi^{\prime}(-\frac{1}{n-1})$ indicates how much the loss decreases if the negative-pair similarity falls below $-\frac{1}{n-1}$, while $\phi^{\prime}(1)$ measures how much the loss increases if the positive-pair similarity drops below $1$. When the condition in (8) is satisfied, ignoring the scaling factor, the loss reduction from separating negative pairs beyond the similarity threshold of $-\frac{1}{n-1}$ can be greater than the loss increase from reducing positive-pair similarity below $1$. As a result, the optimization process favors pushing negative pairs even further apart, leading to the excessive separation. We provide a specific loss where this issue arises:
@@
Mitigating Excessive Separation.
A natural question that arises is whether the excessive separation of negative pairs in full-batch CL can be mitigated. According to Theorem 5.1 and Theorem 5.3, this issue depends on the specific form of the contrastive loss. In particular, under the independently additive loss $\mathcal{L}_{\operatorname{ind-add}}({\bm{U}},{\bm{V}})$, the optimal embeddings do not suffer from excessive separation when $c_{2}=1$ (i.e., case (ii) of Theorem 5.1), whereas the issue arises when $c_{2}=0$ and condition (8) holds.
This observation suggests two potential solutions to the excessive separation problem. The first is to set $c_{2}=1$ in the loss, effectively incorporating within-view negative pairs. The second is to tune hyperparameters such that condition (8) does not hold, i.e., case (iii) in Theorem 5.1. While both approaches are theoretically valid, the former offers a more principled and practical remedy. In contrast, the latter lacks clear guidance for selecting suitable hyperparameters and may incur significant computational overhead. Therefore, we advocate including within-view negative pairs as a straightforward and effective strategy to avoid excessive separation in full-batch CL.
@@ 5.2 Mini-Batch Contrastive Learning
In Sec. 5.1, we show that when a proper contrastive loss is chosen for full-batch settings, the learned embedding satisfies the following behavior: all negative pairs have the cosine similarity of $-\frac{1}{n-1}$ and all positive pairs are perfectly aligned. What about the practical scenarios when we use mini-batches for training? Suppose $n$ training samples are partitioned into $b$ mini-batches where each batch contains $m:=n/b$ samples. Consider training the embeddings by under the fixed mini-batch configuration, where $k$-th mini-batch contains the samples with indices in ${\bm{I}}_{k}=\{m(k-1)+1,m(k-1)+2,\cdots,mk\}$. Under this scenario, the below theorem analyzes the behavior of embeddings learned by mini-batch training.
@@
Suppose $d\geq m-1$. Let the contrastive loss $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ be one of the forms in Theorem 5.1. Define $f^{\star}_{\operatorname{batch}}$ as the optimal encoder that minimizes the fixed mini-batch loss, given by
@@
According to Theorem 5.5, the embeddings learned by mini-batch settings have the following behaviors. First, the positive-pair similarity is equal to 1, i.e., all positive pairs are fully aligned. Second, the expectation of negative-pair similarities is equal to $-\frac{1}{n-1}$, which happens for full-batch settings as well. Third, unlike full-batch settings, the negative-pair similarity is not uniform across the pairs, i.e., the variance is positive when the mini-batch size $m$ is strictly less than the sample size $n$. Thus, the effect of using mini-batch (compared with using full-batch) is in the increased variance of negative-pair similarities. Throughout the paper, we call such phenomenon as the excessive separation of negative pairs in mini-batch CL.
Figure 2 visualizes the effect of using mini-batches, compared with full-batch settings, when $n=4$ and $m=2$. For full-batch settings, shown in (c) of Figure 2, the variance of negative-pair similarities is zero, indicating that all negative pairs are equi-distant. In contrast, (a) and (b) of Figure 2 illustrate the embeddings learned by mini-batch settings, where the fixed mini-batches are specified as $\left({\bm{U}}_{[1:2]},{\bm{V}}_{[1:2]}\right)$ and $\left({\bm{U}}_{[3:4]},{\bm{V}}_{[3:4]}\right)$. For both (a) and (b), the variance of negative-pair similarities is positive, where (a) represents the case that achieves the highest variance in (11), and (b) corresponds to the case that achieves the lowest variance.
@@
Note that the variance of negative-pair similarities in (11) depends on the batch size $m$. For example, in the full-batch case where $m=n$, the upper bound in  (11) is zero, which is consistent with Theorem 5.1. One can confirm that the upper and lower bounds on the variance is a monotonically decreasing function of $m$, which implies that smaller batch sizes inherently exacerbates the excessive separation of negative pairs in mini-batch settings.
The below theorem analyzes the effect of batch size on the training dynamics, when a popular CL loss is used:
@@
Proposed Variance Reduction Method.
Prior empirical studies on CL have shown that mini-batch settings often underperform compared to full-batch settings (Chen et al., 2020; Radford et al., 2021). This naturally raises a question: What is the main factor contributing to this performance degradation in mini-batch settings, and how can it be addressed? According to Theorem 5.5, from the perspective of cosine similarity of embeddings, the key difference introduced by mini-batch settings lies in the increased variance of negative-pair similarities. Motivated by this theoretical insight, we propose an approach to improve mini-batch contrastive learning by introducing an auxiliary loss term, $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$, which explicitly reduces the variance of negative-pair similarities:
@@
as the auxiliary loss for reducing the variance of negative-pair similarities.
One can combine arbitrary conventional mini-batch loss $\mathcal{L}\left({\bm{U}}_{[m]},{\bm{V}}_{[m]}\right)$ with the proposed auxiliary loss to get the modified loss, given by
@@
where $\lambda>0$ is a hyperparameter. By including the proposed term into the mini-batch loss, we encourage all negative-pair similarities to be close to $-\frac{1}{n-1}$, the ideal value achieved in full-batch settings in Theorem 5.1. As a result, the proposed loss controls the variance of negative-pair similarities.
@@ 6 Empirical Validation
In this section, we empirically validate the impact of our theoretical results discussed in Sec. 5, especially for the practical scenarios of mini-batch settings. First, we empirically observe that the excessive separation of negative pairs (proven in Theorem 5.5) actually happens in experiments on benchmark datasets. Second, we empirically confirm that such excessive separation issue can be mitigated by using the proposed loss in Def. 5.7 which reduces the variance of the negative-pair similarities. Third, we observe such variance reduction improves the quality of learned representations in various real-world experiments.
@@ 6.1 Excessive Separation of Negative Pairs
To evaluate the effectiveness of our proposed auxiliary loss, we train five models for each batch size by minimizing the SimCLR loss combined with the auxiliary loss $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$ in Def. 5.7 with the hyperparameter of $\lambda=30$. The variances of negative-pair similarities from these additional models are shown in the last column of Table 1, and are consistently reduced across all batch sizes. This indicates that our proposed loss effectively mitigates excessive separation of negative pairs in mini-batch settings.
@@ 6.2 Effect of Variance Reduction on Performance
We further investigate whether reducing the variance of negative-pair similarities improves the quality of learned representations in terms of the downstream performances.
@@
We pretrain models on CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009) using various contrastive losses that follow the formulation in Def. 3.1, including SimCLR, DCL, and DHEL. For all methods, we compare models trained with and without incorporating the auxiliary loss $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$ in Def. 5.7. The hyperparameter $\lambda$ for the proposed loss is tuned over $\{0.1,0.3,1,3,10,30,100\}$. Unless otherwise specified, the other settings follow those in Sec. 6.1.
@@
Figure 3: Classification accuracy on CIFAR datasets. Models are trained by minimizing the SimCLR loss with and without the auxiliary loss proposed in Def. 5.7, using various temperature parameters in the SimCLR loss.
The quality of the representations learned through CL is known to be sensitive to the choice of the temperature parameter in contrastive losses, as it influences the distribution of similarities among embeddings (Wang & Liu, 2021). To investigate this sensitivity, we train models using the SimCLR loss with temperature values ranging from 0.07 to 2.00. We compare the standard SimCLR loss against our proposed variant, which incorporates the auxiliary loss introduced in Def. 5.7 to reduce the variance of negative-pair similarities. As shown in Figure 3, incorporating the proposed term leads to a consistently higher and more stable classification accuracy across all temperature settings, alleviating the need for careful temperature tuning.
Figure 4: Effect of the auxiliary loss proposed in Def. 5.7 on top-1 classification accuracy when combined with various baseline methods (SimCLR, DCL, and DHEL) on CIFAR datasets. The gray bars highlight the performance gains achieved by incorporating the proposed term across various batch sizes. The proposed auxiliary loss consistently improves the model performance.
We further evaluate the auxiliary loss in Def. 5.7 on existing CL methods, including DCL and DHEL. Experiments are conducted on the CIFAR datasets with various batch sizes. As shown in Figure 4, incorporating the proposed term leads to improved classification accuracy, with the effect being more pronounced at smaller batch sizes. Additional results on the ImageNet dataset are presented in Appendix E.